OpenStudy (anonymous):

help! the problem is: write a quadratic function in standard form with zeros 6 and -8 (the answer is f(x)=x^2+2x-48)

jimthompson5910 (jim_thompson5910):

x=6 or x=-8 x-6=0 or x+8=0 (x-6)(x+8)=0 x^2+8x-6x-48=0 x^2+2x-48=0 So the function f(x)=x^2+2x-48 has the zeros 6 and -8